OpenStudy (anonymous):

Differential Equations: a) Find the general solution of differential equation y''+y'-6y=0 b) Find the solution of the same equation satisfying the initial condition y(0)=1, y'(0)= a. Then find a so that the solution approaches zero as t goes to + infinity.

OpenStudy (amistre64):

constant coeffs are solved with an equivalent quadratic

OpenStudy (amistre64):

r^2 +3 -6 = 0 y = c1 e^{r1x} + c2 e^{r2x}

OpenStudy (lalaly):

let y=e^(rx) \[r^2+r-6=0\]\[(r+3)(r-2)=0\]\[r=-3,2\]\[y=c_1e^{-3x}+c_2e^{2x}\]

OpenStudy (amistre64):

that 3 spose to be an r lol

OpenStudy (amistre64):

part b is simply take the derivative of "y" and to solve for the boundary value stuff

OpenStudy (amistre64):

i got no idea how that t to infinity part fits in tho

OpenStudy (anonymous):

that's cool. I'll try to figure that out. Thanks all!

OpenStudy (amistre64):

good luck, if I could see a"t" that could be limited that would help; but i think that is more material based to what you have and we lack.